Standing on a balcony, you throw your keys to a friend standing on the ground below. 1.2 seconds after you release the keys, they have an instantaneous velocity of 14.0 m/s, directed 58° below the horizontal. What initial velocity did you give them?

magnitude
direction below the horizon
Answer #1

here,

the time taken , t1 = 1.2 s

final vertical speed , v = 14 m/s

theta = 58 degree

for vertical direction

using first equation of motion

v * sin(theta) = uy + g * t

14 * sin(58 ) = uy + 9.81 * 1.2

uy = 0 m/s

for horizontal direction

the initial horizontal velocity , ux = final horizontal velocity

ux = v * cos(theta)

ux = 14 * cos(58) = 7.42 m/s

the initial speed of key us 7.42 m/s and 0 degre below the horizontal